# Conditional Quantile Estimation for Uncertain Watch Time in Short-Video Recommendation

### v1

@@ 5. Conclusion
In this paper, we introduce CQE (Conditional Quantile Estimation), a novel approach to modeling the conditional distribution of uncertain watch time in short video recommendation systems. By leveraging quantile regression, CQE captures the inherent uncertainty and variability in user behavior, providing a more accurate and robust estimation of user engagement. Furthermore, we introduce innovative strategies for integrating the CQE model into existing recommendation systems, including CDE (Conditional Expectation), CSE (Conservative Estimation), and DQC (Dynamic Quantile Combination). These strategies leverage the estimated watch time distribution to enhance user satisfaction and engagement through personalized recommendations. Through extensive offline evaluations on public datasets and a real-world deployment on a platform with over 300 million daily active users, we demonstrated the superior performance of CQE compared to state-of-the-art methods.

### v4

@@ 5. Conclusion
In this paper, we propose the Conditional Quantile Estimation (CQE) framework to model the uncertainty in watch-time prediction for short-video recommendation. Unlike traditional methods that estimate a single expected watch time, CQE captures the entire conditional distribution using quantile regression, allowing for more flexible and personalized recommendations. We further introduce three recommendation strategies—Conservative Estimation (CSE), Dynamic Quantile Combination (DQC), and Conditional Expectation (CDE)—tailored for different scenarios. Extensive offline experiments and large-scale online A/B testing demonstrate CQE’s effectiveness, leading to significant improvements in user engagement metrics such as active days and watch time.
For future work, we aim to explore dynamic quantile combination with the whole range of quantiles, integrate multi-modal features to improve content understanding, and investigate reinforcement learning-based optimization for automated strategy selection. Further research can also focus on bias mitigation to enhance fairness in recommendation.
Appendix A Supplementary Results
Similar to our findings with DeepFM and AutoInt (Table 5), experiments with DCNV2 (Wang et al., 2021) backbone (Table 7) demonstrate CQE’s consistent superior performance in user interest modeling.
Table 7. Comparison between CQE and other methods in user interest prediction task with backbone DCNV2. Boldface means the best-performed methods.
Backbone Method KuaiRand WeChat
GAUC nDCG@1 nDCG@3 nDCG@5 GAUC nDCG@1 nDCG@3 nDCG@5
DCNV2 PCR + MSE 0.601 0.428 0.481 0.510 0.542 0.569 0.571 0.577
PCR + CE 0.638 0.435 0.486 0.519 0.555 0.588 0.589 0.599
PCR + CQECDE(Ours) 0.643 0.437 0.492 0.523 0.595 0.654 0.650 0.651
WTG + MSE 0.601 0.418 0.469 0.508 0.539 0.571 0.577 0.587
WTG + CE 0.649 0.443 0.495 0.528 0.558 0.560 0.569 0.579
WTG + CQECDE(Ours) 0.652 0.449 0.497 0.530 0.586 0.637 0.638 0.640
$\text{D}^{2}$o + MSE 0.620 0.431 0.481 0.525 0.525 0.561 0.568 0.571
$\text{D}^{2}$o + CE 0.659 0.448 0.498 0.531 0.534 0.571 0.579 0.592
$\text{D}^{2}$o + CQECDE(Ours) 0.663 0.455 0.502 0.533 0.565 0.638 0.621 0.625
Appendix B Supplementary Code
We provide the implementation of the Pinball Loss function of Equation 4 in both PyTorch and TensorFlow.
B.1. PyTorch Implementation
1import torch
2
3def pinball_loss(y_true, y_pred, tau=0.5):
4 error = y_true - y_pred
5 return torch.mean(torch.maximum(tau * error, (tau - 1) * error))
Listing 1: Pinball Loss in PyTorch
B.2. TensorFlow Implementation
1import tensorflow as tf
2
3def pinball_loss(y_true, y_pred, tau=0.5):
4 error = y_true - y_pred
5 return tf.reduce_mean(tf.maximum(tau * error, (tau - 1) * error))
Listing 2: Pinball Loss in TensorFlow
This loss function is essential for quantile regression and serves as a core component of our CQE model.